# Generic motives and motivic cohomology of fields

### v1

@@ Abstract.
These notes are devoted to the study of generic motives and their application to the structure of the motivic cohomology of fields. Related to Voevodsky’s homotopy $t$-structure on motivic complexes, generic motives are also central to the conjectural motivic $t$-structure, as was shown by Beilinson. Morphisms of generic motives are closely related to the functoriality of Rost’s cycle modules. We obtain some new computations of generic motives, and deduce computations of morphisms of generic motives, which draw a conjectural picture.
The structure of generic motives reflects into that of motivic cohomology of fields. After giving a new argument for determining weights in the K-theory of number fields using Borel’s isomorphism, we deduce that many motivic cohomology groups of fields have infinite rank, usually of the same cardinality as the field itself. As an example motivic cohomology groups of the real or complex numbers are shown to be uncountable in many degrees and twists. Despite these results, we formulate a hope on their behavior with respect to transcendental elements, providing a counterpart to the Beilinson-Soulé vanishing conjecture.
This paper is dedicated to the memory of Jacob Murre, whose enthusiasm for, and commitment to, Grothendieck’s theory of motives was a driving force for me.
@@ 1. Introduction
One of the acts of birth of motivic cohomology was, retrospectively, Borel’s 1974 computation of the K-theory of number fields as a corollary of that of the stable real cohomology of arithmetic groups [Bor74]. Higher K-theory had barely been introduced by Quillen in his celebrated paper [Qui73b], and Lichtenbaum had just embarked in a lifetime quest to develop higher regulators and analogs of Dirichlet’s class number formula. Both works were published in the same lecture notes in 1973. The next step seemed to have been crystallized by two works in the year 1979: the first one is a paper due to Deligne [Del79] on special values of L-functions attached to motives, and the second one consists of lectures of Spencer Bloch, on algebraic cycles. They both lead to the second age of motives, led by the famous conjectures due to Beilinson, stated in [Bi84] in 1984. Not long after the statement of these conjecture, Voevodsky’s PhD proposed a new approach via the famous h-topology, and the category of h-motivic complexes. This opened the way to a totally new era, which encompassed mixed motivic complexes into a homotopy theory for algebraic variety, now known as motivic homotopy theory. It led to many success, and most notably the solution to the Milnor and Bloch-Kato conjectures by Voevodsky ([Voe11]), with important contributions by Rost ([Ros96]).
At the time being, Beilinson’s conjectures on motivic complexes have been positively answered, with the construction of several suitable categories of mixed motives, equipped with six functors and an $\ell$-adic realization functor compatible with the 6 operations (see [Voe92, VSF00, Ayo07, CD19, Ayo14, CD16, CD15, Par21, Spi18]). [^1] The publication of [CD19], available on arXiv in its almost final form since 2009, has been delayed more than reasonable by unfortunate events independent of the will of the authors. One main problem remains open, the question of the motivic t-structure. Though its definition is clear, its properties are bond to conjectures, such as the Beilinson-Soulé conjecture, which can be reduced to the question of the motivic cohomology of fields. This is one of the reason to consider the so-called generic motives which is one of the central notion adressed in these notes.
Recall also that Voevodsky’s theory of motivic complexes $\operatorname{DM}(k)$ over a perfect field $k$ comes with a natural $t$-structure, called the homotopy $t$-structure. Generic motives were introduced in [Dég02] in order to study the objects of its heart, the homotopy modules with transfers. [^2] Recall on Voevodsky’s theory, effective and stable, and the homotopy $t$-structure are given in Section 2.1. One of the main results of Voevodsky is that the corresponding sheaves (homotopy invariant sheaves with transfers) admits a Gersten resolution. Generic motives allow to enhance this result, showing that Voevodsky’s homotopy modules with transfers are equivalent to Rost’s cycle modules (see 3.3.7 for recall). Indeed, one of the main part of this result is to be able to recover the rich functoriality of cycle modules out of that of homotopy modules. One of the result of [Dég02] is that this functoriality is completely encoded into morphisms of generic motives (see recall in Section 3.3). The first goal of these notes is to improve the theory of generic motives, in several direction.
It is time to recall that the generic motive of a function field333finitely generated field extension $F$ over $k$ is the pro-motive associated with all possible smooth schematic models $X$ over $k$:
@@
where $M(X)$ is the homological motive over $k$ associated with $X$ — see also 3.1.4. Such a generic motive can be twisted $\operatorname{\hat{M}}(F)\{n\}$ with respect to an arbitray integer $n\in\mathbb{Z}$, with respect to $\mathbb{G}_{m}$-twists: $\mathbbm{1}\{1\}:=\mathbbm{1}(1)[1]\subset M(\mathbb{G}_{m})$.
Fundamentally, generic motives form a conservative family of motivic points, with respect to the homotopy $t$-structure on $\operatorname{DM}(k)$. In other words, $\operatorname{\hat{M}}(F)\{i\}$ is derived pro-$t$-projective with respect to the $t$-category $\operatorname{DM}(k)$: see 3.2.10. This echoes Beilinson’s notion of correspondences at the generic point, which were introduced in [Bei02]. In his work, Beilinson considers the dual of generic motives, which are ind-cohomological motives; see 3.2.4 for details. In fact, he also proves that the latter allows to conjecturally build a conservative of derived $t$-injective objects with respect to the motivic $t$-structure on the triangulated category of $n$-motives, for an arbitrary large integer $n\geq 0$; see [Bei02, Prop. 4.5] for details.
Therefore, the category of generic motives is fundamental to understand motivic complexes. In particular, its morphisms can be considered as specialization morphisms, and admits a (partial) description in terms of the functoriality of cycle modules. We improve the computations of morphisms by explicitly computing generic motives associated with low dimensional varieties: curves and surfaces. This illustratesd Beilinson’s theory as it appears that the basic bricks of motives, given by jacobians for curves, and by the so-called transcendental motives of Kahn, Pedrini and Murre, appears as constituent of the corresponding generic motives of their function fields:
@@
    If $X=C$ is a curve with Albanese scheme $A$ (dual of the Jacobian), there exists a homotopy exact sequence of pro-motives over $k$:
    $\underset{x\in C_{(0)}}{\text{\rm``$\prod$''}}\operatorname{\hat{M}}(\kappa_{x})\{1\}\xrightarrow{\prod_{x}\partial_{x}}\operatorname{\hat{M}}(K)\xrightarrow{(\pi_{0}^{K},\pi_{1}^{K})}\mathbbm{1}\oplus\underline{A}$
    which is split if and only if $C$ is rational. We refer to 4.2.2 for more details and notation.
@@
    If $X=S$ is a surface, with Albanese scheme $A$, first Betti number $b_{1}$, and homological transcendental motive $\operatorname{M}_{2}^{tr}(X)$, one gets two homotopy exact sequences of pro-motives:
    ${\underset{s\in S^{(2)}}{\text{\rm``$\prod$''}}\operatorname{\hat{M}}(\kappa_{s})\{2\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\underset{x\in S^{(1)}}{\text{\rm``$\prod$''}}\operatorname{\hat{M}}(\kappa_{x})\{1\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\prod_{x}\partial_{x}}$^(${\operatorname{\hat{M}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\hat{M}}(S^{\leq 1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\mathbbm{1}\oplus\underline{A}\oplus b_{1}.\mathbbm{1}(1)[2]\oplus\operatorname{M}_{2}^{tr}(S)\oplus\underline{A}(1)[2]\oplus\mathbbm{1}(2)[4]$
    The pro-motive $\operatorname{M}(S^{\leq 1})$ has a precise interpretation in term of the coniveau filtration: see 4.3.7 for details.
Generic motives are therefore extensions (non-split in general) of the fundamental bricks of (mixed) motives. Both results shed new light on morphisms of generic motives. They confirm the fact that morphisms of generic motives should be topologically generated by Rost’s four basic maps (see Conjecture).
Generic motives reflect the properties of motivic cohomology of fields. As stated previously, determining the latter is the next fundamental questions one faces in Beilinson’s approach to mixed motives. At the time being, the only results at our disposal are: Quillen’s finiteness theorem for the K-theory of number rings [Qui73a] and computation of algebraic K-theory of finite field [Qui72], Borel’s computation of the rank of algebraic K-theory of number rings [Bor74], and Harder’s analogous computation for function fields of curves over a finite field [Har77].
It is known how to compute the rank of motivic cohomology of number fields and rings (see the precise formulation in 5.1.1), out of the computation of Borel. It consists in comparing Borel and Beilinson’s regulators: see [BG02]. This is rather indirect. By studying the compatibility of Borel’s computation with $\lambda$-filtrations at each step, we give a direct proof of this computation in Section 5.1.
The main result of these notes, deduced from the computation of generic motives and the preceding result of Borel is the fact motivic cohomology $\operatorname{H}_{\mathrm{M}}^{**}(K)$ of a field $K$ of characteristic 00 is infinite in a wide range. More precisely, we prove:
@@
Let $K$ be a field of characteristic 00 of transcendence degree $d>0$.
Then for any $n\in[2,d+1]$ and any $i\geq n$, the abelian group $\operatorname{H}_{\mathrm{M}}^{n,i}(K)$ has infinite rank equal to $\operatorname{card}(K)$.
In particular, motivic cohomology of the real or complex numbers is uncountable in the indicated range. We refer the reader respectively to 5.2.2 and 5.2.4 in the text for proofs. In the range $n=i$, where motivic cohomology agrees with Milnor K-theory, the result was already known from Bass and Tate [BT73, Prop. 5.10]. For further information, in particular in positive characteristic, we refer the reader to Section Section 5.2.
Though this theorem shows that motivic cohomology of fields is usually enormous, the behavior with respect to purely transcendental extensions suggests the following plausible conjecture:
Conjecture C.
Let $K$ be a field finitely generated over its prime field $F$, with Krull dimension:
$\delta(K)=\begin{cases}1+\operatorname{degtr}(K/\mathbb{Q})&F=\mathbb{Q}\\ \operatorname{degtr}(K/\mathbb{F}_{p})&F=\mathbb{F}_{p}\end{cases}$
Then, one has the following vanishing of the rational motivic cohomology of $K$
@@
The conjecture is known in any case for $\delta(K)\leq 1$. We refer the reader to the end of Section Section 5.2 for further discussion.

### v3

@@ Abstract.
This paper investigates the structure of generic motives and their implications for the motivic cohomology of fields. Originating in Voevodsky’s theory of motives and related to Beilinson’s vision of a motivic $t$-structure, generic motives serve as pro-objects encoding essential information about cycles and cohomology. We present new computations of generic motives, focusing on curves and surfaces. These computations suggest a conjectural framework for morphisms of generic motives and highlight the central role of transcendental motives.
We then focus on the motivic cohomology of fields, building on Borel’s rank computation of K-theory and its relation to higher regulators. We provide a direct argument for determining the weights in the $\lambda$-structure of the K-theory of number fields, bypassing the need for regulator maps. We show that motivic cohomology groups are often of infinite rank, typically matching the cardinality of the base field. For instance, we prove that motivic cohomology groups of $\mathbb{R}$ and $\mathbb{C}$ are uncountable in many bi-degrees. Despite this, we propose a conjecture that complements the Beilinson-Soulé vanishing conjecture, suggesting that the growth of motivic cohomology is more controlled than these results may initially indicate.
This paper is dedicated to the memory of Jacob Murre.
@@ 1. Introduction
1.1. Historical perspective
Beilinson’s conjectures
The concept of motivic cohomology was first formulated, as it seems, by Beilinson in a 1982 letter to Soulé [^1] This letter was once available on the K-theory preprint server, which unfortunately went offline. The first published reference by Beilinson mentioning motivic cohomology (under the name “absolute” cohomology) is [Bei84]. as a universal cohomology theory for algebraic varieties, analogous to singular cohomology in topology. Inspired initially by Bloch’s construction of higher regulators for $K_{2}$, Beilinson envisioned motivic cohomology as a universal receptacle for characteristic classes, serving as a refinement of Quillen’s higher algebraic K-theory.
This vision was preceded by a series of foundational computations linking $K$-theory and special values of zeta functions. Early results by Moore, Tate, and Garland on the group $K_{2}$ of global fields (see Bass’s Bourbaki talk [Bas71]) led to the formulation of the famous Birch-Tate conjecture — much progress has been made on this conjecture, but the general case remains open. This motivated Lichtenbaum to propose a series of synthetic conjectures, first in the context of the zeta function of a totally real number field, linking special values, $K$-theory, étale cohomology, and higher regulators (see [Lic73]). A decisive breakthrough came with Borel’s results in 1972, which extended Garland’s earlier work and determined the ranks of higher algebraic $K$-groups of rings of integers in number fields; see [Bor72, Bor74]. Borel also established the existence of higher regulators in this setting and verified Lichtenbaum’s conjecture in the case of (arbitrary) number fields; see [Bor77, Bor80].
These converging developments culminated in a profound conceptual synthesis through the work of Beilinson, who brought together several deep lines of thought into a unified vision. Bloch’s 1978 Irvine lectures (published later in [Blo00]), in which he revisited Borel’s work on regulators of number fields through the lens of dilogarithms and proposed a visionary extension to the case of elliptic curves, had a decisive influence. Around the same time, Deligne’s study of special values of $L$-functions [Del79] brought Grothendieck’s theory of motives into the picture. Building on these inputs, Beilinson formulated a striking and coherent picture: combining the conjectural theory of mixed motives with Deligne’s theory of mixed Hodge structures, he expressed the higher regulator as a byproduct of the realization functor from mixed motives to mixed Hodge structures.
Returning to the universal property of motivic cohomology, Beilinson’s regulator also admits a concrete, unconditional formulation: rational motivic cohomology can be defined using the $\lambda$-structure on algebraic $K$-theory as developed by Soulé in [Sou85]. The Beilinson regulator then appears as the universal map to Deligne cohomology, induced by Chern classes. This allowed Beilinson to state conjectures predicting the special values of $L$-functions of motives in terms of this regulator, encompassing the earlier predictions of Lichtenbaum, and including, up to rational factors, both the Birch–Tate and Birch-Swinnerton-Dyer conjectures; see [Bei84, Bei87].
Beilinson’s work proved extremely fruitful and inspired many mathematicians to approach his conjectures from a variety of angles. Bloch was the first to propose in [Blo00] an integral version of motivic cohomology through the theory of higher Chow groups, providing a concrete model compatible with Beilinson’s conjectures. Suslin slightly later introduced in [SV96] an alternative approach via Suslin homology, inspired by topological methods.
Voevodsky’s motivic theory
It was Voevodsky who brought a decisive and revolutionary shift. In a groundbreaking PhD thesis, he introduced a new homotopical viewpoint based on the contractibility of the affine line and the use of the h-topology. Building on this approach, he developed a new framework for defining Beilinson’s conjectural motivic complexes that not only met the formal expectations envisioned by Beilinson, but also integrated naturally into a broader theory of motivic homotopy, later developed in collaboration with Morel. This framework led to several landmark results, including the proofs of the Milnor and Bloch–Kato conjectures [Voe11], as well as the Quillen–Lichtenbaum conjecture. Voevodsky’s approach laid the foundations for a complete theory of motivic complexes, modeled on Grothendieck’s $\ell$-adic formalism, as anticipated by Beilinson. It was brought to maturity through the development of a full Grothendieck six-functor formalism, both in its Nisnevich and étale variants (see [Voe92, VSF00, Ayo07a, Voe10b, RØ08, Ivo11, CD19, Ayo14, CD16, CD15, Par21, Spi18]). [^2] We review the developments of the Nisnevich variant in 2.1.6.
A concise account of the construction of the category of motivic complexes $\operatorname{DM}(k)$ over a (perfect) field will be provided in Section 2.1. For the purposes of this work, we have chosen to present this construction within the framework of $\infty$-categories, which, for many applications in homological algebra, are better suited than triangulated categories. To our knowledge, this is the first presentation of motivic complexes directly formulated in terms of $\infty$-categories. To assist readers unfamiliar with these concepts, we have included a detailed overview on $\infty$-categories in the appendix 6.333The $\infty$-category of motivic complexes is denoted by $\operatorname{\mathscr{D}M}(k)$ while its homotopy category, equivalent to the non-effective version of Voevodsky’s construction, is denoted by $\operatorname{DM}(k)$.
1.2. Content of the paper
Motivic cohomology of fields
Despite these advances, one major problem remains open: the motivic $t$-structure whose heart would provide the abelian category of mixed motives. With rational coefficients and other a base field, Beilinson’s conjectural description leaves no choice, as it demands that the $\ell$-realization is conservative and $t$-exact, with respect to the canonical $t$-structure on the derived category of $\ell$-adic étale sheaves (Galois modules in that case). Therefore a rational geometric motivic complex should be non-negative (resp. positive) for the motivic $t$-structure if and only if its $\ell$-adic realization is. However, the fact that this actually defines a $t$-structure [^4] Orthogonality and existence of truncations are problematic depends on at least two deep conjectures: the conservativity conjecture (see [Ayo17]) and the Beilinson-Soulé vanishing conjecture. In fact, Levine showed that the latter is equivalent to the fact that the above definition induces a $t$-structure on the sub-category of rational geometric Artin-Tate motives; see [Lev93].
The Beilinson-Soulé conjecture asserts that for any smooth scheme $X$ over a field $k$, rational motivic cohomology
$\operatorname{H}_{\mathcyr{B}}^{n,i}(X):=\operatorname{Gr}_{\gamma}^{i}K_{2i-n}(X)_{\mathbb{Q}}$
vanishes if $n\leq 0$ and $(n,i)\neq(0,0)$ (see [BMS87, Introduction]). Using the coniveau spectral sequence, this conjecture can be reduced to the case of arbitrary function fields. [^5] More precisely, the residue fields of all (schematic) points of $X$. In particular, the key issue becomes our understanding of the motivic cohomology of fields.
For now, rational motivic cohomology of fields is known for finite fields and global fields, in other words, fields $K$ of Kronecker dimension less or equal to one, where one defines this degree as:
$\delta(K)=\begin{cases}1+\operatorname{trdeg}(K/\mathbb{Q})&\text{if }\operatorname{char}(F)=0,\\ \operatorname{trdeg}(K/\mathbb{F}_{p})&\text{if }\operatorname{char}(F)=p.\end{cases}$
Indeed, in positive characteristic $p$, if $\delta(K)\leq 1$, results of Quillen (finite fields) and Harder imply that:
$\operatorname{H}_{\mathcyr{B}}^{n,i}(K)=\begin{cases}\mathbb{Q}&n=i=0,\\ K^{\times}\otimes\mathbb{Q}&n=i=1,\\ 0&\text{otherwise.}\end{cases}$
The case of number fields.
This is more subtle than the positive characteristic case. As mentioned in the preceding historical review, Borel computed the ranks of the rational K-groups of $K$, which, in degrees $>1$, coincide with those of its ring of integers $\mathcal{O}_{K}$. These groups are concentrated in odd degrees $12n-$ 1. Borel also defined a regulator map: [^6] See (5.1.9.a) below for an explicit definition.
$\rho_{Bo}:K_{2n-1}(\mathcal{O}_{K})_{\mathbb{Q}}\rightarrow V_{n}$
with values in a finite-dimensional real vector space $V_{n}$, and showed it is injective. On the other hand, Beilinson defined a regulator map, as mentioned above, with values in the Deligne-Beilinson cohomology of $K$
$\rho_{Be}:\operatorname{H}_{\mathcyr{B}}^{n,i}(K)\rightarrow H_{\mathcal{D}}^{n}(\operatorname{\mathrm{Spec}}(K),\mathbb{R}(i))$
and claimed that this map agrees with Borel’s regulator, once one takes into account the Chern character isomorphism:
$K_{i}(X)_{\mathbb{Q}}\simeq\oplus_{n\geq 0}\operatorname{H}_{\mathcyr{B}}^{2n-i,n}(X).$
As a consequence of the injectivity of Borel’s regulator, one deduces that:
(1.2.vanish) $\forall(n,i)\in\mathbb{Z}^{2},\operatorname{H}_{\mathcyr{B}}^{n,i}(K)=0\text{ unless }n=i=0\text{ or }n=1.$
This, in particular, implies the Beilinson-Soulé conjecture for number fields.
The first written comparison between Borel’s and Beilinson’s regulators was given by Rapoport in [Rap88]. Burgos later provided a more complete and accurate treatment in [BG02], and corrected a discrepancy by identifying a missing factor of two in the original comparison.
In particular, the whole procedure to deduce the computation of motivic cohomology of number fields is intricate and technically involved. The first task we undertake in this work is to give a direct proof of (1.2.vanish), without appealing to any regulator map. We approach this result through a direct analysis of Borel’s isomorphism, which uses the relation of the K-theory of $\mathcal{O}_{K}$ with the stable real cohomology of the arithmetic group $\operatorname{SL}_{r}(\mathcal{O}_{K})$. Our strategy is to examine the compatibility of each step in Borel’s construction with natural $\lambda$-ring structures, starting from the canonical one on $K_{n}(\mathcal{O}_{K})\otimes_{\mathbb{Q}}\mathbb{R}$. This ultimately reduces the proof of (1.2.vanish) to determining the $\lambda$-weights of the indecomposable part of the real homology of the “compact twin” of the real Lie group $\operatorname{SL}_{r}(K\otimes_{\mathbb{Q}}\mathbb{R})$.
We refer the reader to 5.1.13 for the final determination of the $\lambda$-weights, and to Section 5.1 for a detailed explanation of the Borel isomorphism and its compatibility with $\lambda$-structures. In this context, the referee pointed out the approach of Borel and Yang on the rank conjecture (see [BY94]), which makes use of partly related techniques. See also [DJ02] for a related result concerning weights in the homology of $\mathrm{GL}_{n}$.
Generic motives
Generic motives were introduced in [Bei02] and in [Dég02] with distinct but complementary objectives. They are defined as pro-objects in the category of motivic complexes $\operatorname{DM}(k)$ over a perfect field $k$, associated with a function field $E$ over $k$ (a finitely generated extension field) and constructed from all possible smooth schematic models $X$ over $k$:
@@ 1.2. Content of the paper
where $M(X)$ denotes the homological motive associated with $X/k$ — see also 3.1.4. Generic motives should be considered with twists $\operatorname{\hat{M}}(E)\{n\}$ for arbitrary integers $n\in\mathbb{Z}$, defined via $\mathbb{G}_{m}$-twists: $\mathbbm{1}\{1\}:=\mathbbm{1}(1)[1]\subset M(\mathbb{G}_{m})$. This definition underscores the need for the $\infty$-categorical framework: in general, the pro-objects of a triangulated category do not themselves form a triangulated category, while pro-objects of an $\infty$-category naturally inherit the structure of an $\infty$-category (see Section 6.6).
In Beilinson’s work, generic motives were employed alongside the filtration by dimension on the category of motivic complexes $\operatorname{DM}(k)$, to characterize uniquely the motivic t-structure, independently of any realization functor. In contrast, in the author’s PhD work, generic motives were used to encode the structure of Rost cycle modules. Specifically, the morphisms of generic motives completely capture the structural functoriality of cycle modules, allowing any motivic complex to give rise to such objects. Applying this result to motivic cohomology, one deduces that motivic cohomology of fields can be organized in a family of cycle modules: fixing an integer $n\in\mathbb{Z}$, one gets $\mathbb{Z}$-graded abelian groups associated with finitely generated extensions $E/k$:
$\operatorname{H}_{\mathrm{M}}^{n+r,r}(E),r\in\mathbb{Z}$
equipped with 4 basic operations: (D1) corestriction maps, (D2) restriction of norm maps, (D3) action of units of $E$ and more generally of Milnor K-theory of $E$, (D4) residues. [^7] The reader is referred to Section 3.3 for more details.
Furthermore, generic motives play a non-conjectural role in relation with the homotopy $t$-structure, analogous to the role envisioned by Beilinson for the conjectural motivic $t$-structure. The existence of the homotopy $t$-structure on the effective motivic complexes $\operatorname{DM^{eff}}(k)$ is a corollary of Voevodsky’s fundamental theorems on homotopy invariant sheaves with transfers, which constitute its heart. This $t$-structure can be extended to the (non-effective or stable) category $\operatorname{DM}(k)$ (We recall this aspect in 2.1.4). In this framework, generic motives function like projective objects with respect to the homotopy $t$-structure on $\operatorname{DM}(k)$, and evaluation at generic motives gives a conservative family. This aspect will be explained in a section recalling the theory of generic motives; see in particular 3.1.14 for the last assertion.
Therefore, the category of generic motives plays a central role in the study of motivic complexes. In particular, the morphisms between generic motives can be interpreted as specialization morphisms, admitting a (partial) description in terms of the functoriality of cycle modules. In this work, we refine the computation of these morphisms by explicitly determining the generic motives associated with low-dimensional varieties, specifically curves and surfaces. This illustrates Beilinson’s considerations on generic motives, as it appears that the fundamental building blocks of motives naturally emerge as constituents of the corresponding generic motives of their function fields. In the case of curves, Jacobians provide the basic bricks, while for surfaces, the so-called transcendental motives of Kahn, Murre, and Pedrini play a similar foundational role.
@@
    If $X=C$ is a curve with Albanese scheme $A$ (dual of the Jacobian), there exists a homotopy exact sequence888here, we use the language of stable $\infty$-categories: the term homotopy exact means both that the first term is the homotopy fiber of the right map and the third term is the homotopy cofiber of the left map. Consequently, the sequence viewed in the associated homotopy category becomes a distinguished triangle. See further 6.4.3. of pro-motives over $k$:
    $\underset{x\in C^{(1)}}{\text{\rm``$\prod$''}}\operatorname{\hat{M}}(\kappa_{x})\{1\}\xrightarrow{\prod_{x}\partial_{x}}\operatorname{\hat{M}}(K)\xrightarrow{(\pi_{0}^{K},\pi_{1}^{K})}\mathbbm{1}\oplus\underline{A}$
    which is split if and only if $C$ is rational. We refer to 4.3.2 for more details and notation.
@@
    If $X=S$ is a surface over a separably closed field, with Albanese scheme $A$, Picard number $\rho$, and homological transcendental motive $\operatorname{M}_{2}^{tr}(X)$, one gets two homotopy exact sequences of pro-motives:
    ${\underset{s\in S^{(2)}}{\text{\rm``$\prod$''}}\operatorname{\hat{M}}(\kappa_{s})\{2\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\underset{x\in S^{(1)}}{\text{\rm``$\prod$''}}\operatorname{\hat{M}}(\kappa_{x})\{1\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\prod_{x}\partial_{x}}$^(${\operatorname{\hat{M}}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\hat{M}}(S^{\leq 1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\mathbbm{1}\oplus\underline{A}\oplus\rho.\mathbbm{1}(1)[2]\oplus\operatorname{M}_{2}^{tr}(S)\oplus\underline{A}(1)[2]\oplus\mathbbm{1}(2)[4]$
    The pro-motive $\operatorname{M}(S^{\leq 1})$ has a precise interpretation in terms of the coniveau filtration: see 4.4.6 for details.
Generic motives are therefore extensions (non-split in general) of the fundamental bricks of (mixed) motives. Both results shed new light on morphisms of generic motives. They confirm the fact that morphisms of generic motives should be topologically generated by Rost’s four basic maps (see 4.2.7).
As a consequence of the preceding result — specifically the first point in the rational case — and of the determination of motivic cohomology of number fields based on Borel’s computation, as previously discussed, we deduce that motivic cohomology of fields is infinite in a wide range of degrees, depending on their Kronecker dimension.
@@
Let $K$ be a field of characteristic 00 of Kronecker dimension $\delta(K)>1$. Then for any $n\in[2,\delta(K)]$ and any $i\geq n$, the abelian group $\operatorname{H}_{\mathrm{M}}^{n,i}(K)$ has infinite rank equal to $\operatorname{card}(K)$.
The proof of this theorem is straightforward, once the generic motive of a pure transcendental extension is understood (4.2.2) and A(1) is established, along with the fact that motivic cohomology groups $\operatorname{H}_{\mathcyr{B}}^{1,n}(E)$ have positive rank whenever $E$ is totally imaginary and $n>1$. The former, obtained as early as 2002 in the author’s PhD thesis [Dég02], generalizes a classical result of Bass and Tate [BT73], while the latter has already been discussed in the historical introduction. However, the preceding consequence for the motivic cohomology of fields, which we believe is significant, seems to have gone unnoticed in the literature so far. We refer the reader to Section 5.2 for further discussion.
Strikingly, motivic cohomology of the real or complex numbers is uncountable in bi-degree $(n,i)$ whenever $i\geq n\geq 2$. In terms of Bloch’s higher Chow groups, this leads to the following result, which, though not strictly an extension of Mumford’s classical geometric theorem on Chow groups [Mum67], shares a similar flavor from an arithmetic perspective, being a consequence of Borel’s profound computation of higher K-groups:
Corollary C.
Let $X$ be a complex smooth algebraic variety, or a real one with a rational point. Then for all integers $(p,q)\in\mathbb{N}$ such that $p\leq q\leq 2p-2$, the higher Chow group:
$\operatorname{CH}^{p}(X,q)$
has infinite, uncountable rank.
It is worth noting that, just as in Mumford’s result for surfaces of general type, the occurrence of infinite rank in our setting is restricted to cycles in codimension $p>1$.
Though the preceding theorem shows that motivic cohomology of fields is usually enormous, the behavior with respect to purely transcendental extensions suggests the following plausible conjecture:
Conjecture D.
Let $K$ be a field finitely generated over its prime field $F$. Then, one has the following vanishing of the rational motivic cohomology of $K$ in terms of its Kronecker dimension $\delta(K)$
@@
The conjecture is known in any case for $\delta(K)\leq 1$. We refer the reader to the end of Section 5.2 for further discussion.